A particle with a net charge of 1.2 mC is placed in a uniform electric field that has a field strength of 5N/C. What is the force on the particle?
Answer #1

force in a electric field qE

q = charge = 1.2mC E = 5N/C

= 1.2*10^-3 * 5 = 6 * 10^-3 N = 6mN